# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v1

@@ 8 Effects of B-cosification and B Values
Figure 5 shows the effects of varying B on the task performance and explanation faithfulness. Classification accuracy initially improves slightly as B increases from 1 to 1.25, benefiting from the extra non-linearity introduced by B>1. However, beyond this point, accuracy declines as higher alignment pressure reduces model flexibility. A similar trend is observed for Comp, peaking around B=1.5 before decreasing. This differs from previous findings in vision models (Böhle et al., 2022), which we attribute to the high sparsity of explanations at larger B values. As alignment pressure increases, fewer tokens receive attribution scores that are not close to zero, leading to poor token importance calibration and lower Comp scores. The effects of B on other metrics are similar and can be found in Appendix L.
@@
Figure 5 also reveals a negative correlation between explanation entropy and B, indicating that higher alignment pressure leads to sparser explanations. This aligns with our expectations: a larger B amplifies the differences between dimensions in $\mathbf{|\text{cos}(x,\hat{W})|^{B-1}}$ of B-cos layers (Equation 2) and the dynamic linear weight assigns more distinct attributions to input features. As a result, explanations become more concentrated, where only a few tokens receive high attributions, while most remain close to zero (cf. Appendix M for an example).
@@
Since B-cos LMs with larger B values rely on fewer tokens for prediction, we investigate whether this may cause them to learn biases in the data. For this, we examine label bias and word-level spurious correlations using the HateXplain dataset, where approximately 60% of training and test examples have positive labels and societal biases are present. Figure 6 shows that a larger B value (B=2.5) reduces the model capacity, leading to a substantially higher prediction positive rate and lower balanced accuracy. Moreover, the B=2.5 model assigns higher attributions to non-semantic [CLS] and [SEP] tokens, indicating a reduced reliance on meaningful content. Notably, this label bias is not observed in the balanced datasets.
@@
We also find that B-cosification—particularly with large B—amplifies the reliance on spurious correlations. For example, the prediction positive rate for examples with the word “black” rises from 49.02% in the test set and 52.94% in the conventional model to 59.80%, 56.86%, and 73.53% in B-cos LMs with B=1, 1.5, and, 2.5, respectively (we provide an example in Appendix N). However, the faithfulness and interpretability of B-cos explanations facilitate the identification of spurious correlations and can effectively guide models toward reducing them (Rao et al., 2023). We leave the exploration of B-cos LMs for bias detection and mitigation to future work.
9 Explanation Efficiency
Beyond improved faithfulness and human interpretability, B-cos explanations are also efficient to extract. Comparing their computational costs with strong post-hoc methods shows that B-cos explanations are the most efficient in both time and memory usage (Table 4).
Method Conv. BERT B-cos BERT
Time (s) Memory (GB) Time (s) Memory (GB)
ShapSampl 37.22 21.95 70.49 22.95
LIME 6.82 21.96 8.92 22.95
SIG 67.46 29.09 108.48 69.32
DecompX 0.76 48.38 - -
B-cos - - 0.08 2.78
Table 4: Computational costs of generating explanations for 100 instances randomly sampled from IMDB (test) using an NVIDIA H100 GPU (batch size 1). We see that the B-cos explanations (bold) are at least 9x faster and require at most $\frac{1}{8}$ of VRAM.
@@ 10 Conclusion
In this work, we introduce B-cos LM, a dynamic linear model that learns task-relevant patterns through increased input-weight alignment pressure. B-cos LMs generate more faithful and human interpretable explanations while maintaining strong task performance and fast convergence. Based on our in-depth analysis of B-cosification, we provide three recommendations for effectively transforming PLMs into B-cos LMs: (1) combine B-cos conversion and task fine-tuning for efficient B-cosification. If resources allow, additional B-cos pre-training can further improve task performance and explanation faithfulness; (2) carefully select the parameter B, as excessively large values can reduce model capacity and lead to overly sparse explanations; and (3) be mindful of biases in training data, especially at high B values, as B-cosification may amplify existing biases.
@@ 11 Limitations
Secondly, our study does not include a direct comparison with other methods designed to enhance model explainability, which may limit the scope of our findings. This omission is due to two reasons: (1) existing explainable models often provide only marginal improvements over post-hoc explanation methods (Brinner and Zarrieß, 2024), and (2) incorporating them into our study would require substantial computational resources, as many baseline explanation methods are computationally expensive.
Finally, although B-cos LMs can be applied to different model architectures and tasks, our experiments focus only on encoder-only models for sequence classification tasks. Extending our approach to other architectures and tasks remains an avenue for future work.
12 Ethical Considerations
@@
All models and datasets used in this work comply with their respective licenses. Their usage aligns with their intended purpose as specified by their creators.
The human study complies with all ethical research guidelines set by our institutes. All participants of the human evaluation study were master’s or doctoral students with backgrounds in computer science or computational linguistics and were proficient in English. They were volunteers and were compensated with the standard hourly salary set by the university (at least 5% above minimum wage). Before participation, all participants were informed about the content and purpose of the study, the collected data and its usage. They were instructed on how they could access, modify, or delete their data post-study and provided their informed consent.
Acknowledgements
This work was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action” - project number 471607914. We also thank all those who provided valuable feedback and suggestions during the development of this work. Their input has been instrumental in improving its quality.
@@ Appendix A Terminology
  • Faithfulness. The extent to which an explanation accurately reflects the model’s actual reasoning process (Jacovi and Goldberg, 2020). A faithful explanation should directly correspond to the internal mechanisms that led to the model’s prediction.
  • Human Interpretability. The ease with which a person can understand the model’s reasoning from the explanation (Lage et al., 2019). A highly interpretable explanation should be clear, concise, and focused on relevant information while avoiding unnecessary or distracting information. However, an explanation that is easy for humans to interpret may not necessarily reflect the model’s actual reasoning process or align with human reasoning patterns.
  • Human Agreement. The degree to which a model’s explanation aligns with the reasoning a human would use for the same prediction. A high-agreement explanation should follow intuitive, logical reasoning patterns similar to human decision-making.
  • Explainability. The extent to which a model’s computations can be faithfully explained and its learned patterns are understandable to humans. A highly explainable model should yield explanations that are both faithful to its actual reasoning process and interpretable to humans.
@@ Appendix B Notation
In this paper, we use lowercase letters for scalars (e.g., b), bold lowercase letters for vectors (e.g., $\mathbf{w}$, $\mathbf{x}$), and bold uppercase letters ($\mathbf{W}$) for matrices. Additionally, we use bold uppercase letters $\mathbf{X}$ and $\mathbf{A}$ to denote a sequence of model inputs or hidden state activations. In § 3, we use $\mathbf{x}$ to denote the input when a function is applied to each element of the input sequence separately. In contrast, we use $\mathbf{X}$ or $\mathbf{A}$ when the function involves interactions between elements, such as in the attention mechanism.
Appendix C Ablation Study
To gain deeper insights into B-cosification, we conduct an ablation study to evaluate the effects of key design choices on model performance. Table 5 reports the effects of these modifications.
Consistent with § 8, B=1 results in worse task performance and lower explanation faithfulness. Using binary cross-entropy (BCE) loss instead of conventional cross-entropy loss has minimal impact on classification accuracy, but leads to better faithfulness results in perturbation-based evaluations. Additionally, architectural adaptations, including removing bias terms and eliminating activation functions in prediction heads, play a crucial role in improving both model performance and explainability in B-cos LMs. Besides, we encountered numerical instability when generating explanations without these architectural adaptations, as the dynamic linear weight for tanh ($\frac{\text{tanh}(\mathbf{x})}{(\mathbf{x})}$) becomes unstable when $\mathbf{x}$ is close to 0.
Beyond ablating components in model design and training, we also examine different explanation methods across models. First, replacing dynamic linear weights $\mathbf{W(x)}$ with gradients for computing input contributions (equivalent to InputXGradient, Kindermans et al., 2016) results in less faithful explanations. Moreover, directly extracting B-cos-like explanations, $\mathbf{W(x)x}$, from a conventional model results in worse faithfulness compared to those from B-cos LMs.. [^4] Extracting $\mathbf{W(x)x}$ from conventional models follows the same approach as in B-cos LMs (cf. § D), except that in standard linear transformations, the dynamic linear weight is replaced by the fixed weight matrix $\mathbf{W}$.
Acc ($\uparrow$) Comp ($\uparrow$) Suff ($\downarrow$) SeqPG ($\uparrow$)
Full system 78.64 59.66 -4.89 77.57
w/o alignment pressure (B=1) 78.07 57.19 -2.57 70.18
w/o BCE training 79.00 49.22 -7.91 79.21
w/o architectural adaptations 77.65 52.23 -3.80 74.30
w/o dynamic linear weights (IxG) 78.64 44.93 -0.60 53.57
$\mathbf{W(x)x}$ from conv. model 80.77 44.92 2.80 70.20
Table 5: Ablation study of key designs in B-cos BERT model on HateXplain. Results are averaged over three runs.
Appendix D Dynamic Linear Representation of Model Components
Here we describe how each model component functions as a dynamic linear module in B-cos LMs.
B-cos Layers
[…]
Non-linear activation functions
In transformer models, non-linearity is typically introduced using (approximately) piecewise linear activation functions, such as ReLU (Nair and Hinton, 2010) and GELU (Hendrycks and Gimpel, 2016). These functions can be easily interpreted as linear transformations with input-dependent weights. For example, $\text{GELU}(\mathbf{x})=\mathbf{x}\times(0.5+0.5\times\text{erf}(\mathbf{x}/\sqrt{2}))$ can be interpreted as a linear transformation where the second term acts as a dynamic linear weight.
Attention block
$\text{Att}(\mathbf{X;Q,}$ $\mathbf{K,V})=\text{softmax}(\mathbf{X^{T}Q^{T}KX)VX}$
$=\mathbf{A(X)VX}=\mathbf{W(X)X}$ (4)
$\text{MSA}\mathbf{(X)=U[W}_{1}\mathbf{(X)X,...,W}_{H}\mathbf{(X)X)]}$ (5)
Appendix E Implementation Details
@@
When constructing examples for SeqPG, we set the sequence length to 50 for AG News, 256 for IMDB, and 25 for HateXplain, aligning with their median lengths. Only examples longer than these thresholds are selected, and they are truncated to construct synthetic examples. Additionally, we only use examples that are correctly predicted with a minimum confidence of 75% after truncation. For a fair comparison, we evaluate B-cos LMs on the same sets of examples constructed based on the predictions of the corresponding conventional models.
Evaluation Setups
For task performance evaluation, we use the complete test set for each task. For faithfulness evaluation, we conduct perturbation-based evaluations on 2000 test examples and SeqPG on 500 test examples for AG News and IMDB. For HateXplain, we use the full test set for perturbation-based evaluation (1,924 examples) and construct 269, 310, and 308 SeqPG examples from it using BERT, DistilBERT, and RoBERTa, respectively.
@@
For B-cos pre-training in § 7, we further pre-train the model on the Wikipedia dataset [^9] https://huggingface.co/datasets/wikimedia/wikipedia using masked language modeling loss with a learning rate of 1e-4 and a 15% masking ratio.
@@
Appendix F SeqPG Example

### v2

@@ 6 Impact of B-cosification and B Values
Figure 5 shows the effects of varying B on the task performance and explanation faithfulness. Classification accuracy initially improves slightly as B increases from 1 to 1.25, benefiting from the extra non-linearity introduced by B>1. However, beyond this point, accuracy declines as higher alignment pressure reduces model flexibility. A similar trend is observed for Comp, which peaks around B=1.5 before decreasing. This differs from previous findings in vision models (Böhle et al., 2022), which we attribute to the high sparsity of explanations at larger B values. As alignment pressure increases, fewer tokens receive attribution scores that are not close to zero, leading to poor token importance calibration and lower Comp scores. The effects of B on other metrics are similar and can be found in Appendix K.
@@
Figure 5 also reveals a negative correlation between explanation entropy and B, indicating that higher alignment pressure leads to sparser explanations. This aligns with our expectations: a larger B amplifies the differences between dimensions in $\mathbf{|\text{cos}(x,\hat{W})|^{\text{B-1}}}$ of B-cos layers (Equation 2) and the dynamic linear weight assigns more distinct attributions to input features. As a result, explanations become more concentrated, where only a few tokens receive high attributions, while most remain close to zero (cf. Appendix L for an example).
@@
Since B-cos LMs with larger B values rely on fewer tokens for prediction, we investigate whether this may cause them to overfit and learn biases in the data. For this, we examine label bias and word-level spurious correlations using HateXplain, where approximately 60% of training and test examples have positive labels and societal biases are present. Figure 6 shows that a larger B value (B=2.5) reduces the model capacity, leading to a substantially higher positive rate in predictions and therefore lower class-balanced accuracy. Moreover, the B=2.5 model assigns higher attributions to non-semantic [CLS] and [SEP] tokens, indicating a reduced reliance on meaningful content. Notably, this label bias is not observed in the balanced AG News and IMDB datasets.
@@
We also find that B-cosification, particularly with large B, amplifies the reliance on spurious correlations. For example, the prediction positive rate for examples with the word “black” rises from 49.02% in the test set and 52.94% in the conventional model to 59.80%, 56.86%, and 73.53% in B-cos LMs with B=1, 1.5, and, 2.5, respectively (we provide an example in Appendix M). However, the faithfulness and interpretability of B-cos explanations facilitate the identification of spurious correlations and can effectively guide models toward reducing them (Rao et al., 2023). We leave the exploration of B-cos LMs for bias detection and mitigation to future work.
7 B-cosifying Decoder-Only Models for Generation Tasks
LLMs are increasingly used as general-purpose assistants, with most based on decoder-only architectures (Zhao et al., 2023; Minaee et al., 2024). While our primary focus is on classification tasks using encoder-only models, we also extend B-cosification to decoder-only models for generation tasks to demonstrate the broader applicability of B-cos LMs. Specifically, we apply B-cosification to two decoder-only models, GPT-2 small (Radford et al., 2019, referred to as GPT-2 afterwards) and Llama-3.2-1B (Dubey et al., 2024, referred to as Llama-3.2 afterwards), and evaluate their language modeling performance and explanation quality on two generation tasks. For more details on the datasets, experimental setup and baseline models, see Appendix C.
Model Probability Gap ($\uparrow$) PPL ($\downarrow$)
BLiMP IOI
GPT-2 0.0055 0.3351 3.10
B-cos GPT-2 0.0059 0.3265 3.04
Llama-3.2 0.0058 0.4652 2.51
B-cos Llama-3.2 0.0065 0.5021 2.64
Table 6: Language ability results for vanilla and B-cos decoder-only models. Probability Gap scores on BLiMP are averaged over nine subsets. Scores where B-cos LM outperforms their vanilla counterparts are in bold. B-cos LMs show language modeling ability comparable to vanilla models. Results for each subset can be found in Table 13 in Appendix N.
B-cosification Setup
Given the complexity of modeling natural language, we use a small B value of 1.1. We do not B-cosify the language head, as its parameters are tied with the embedding layer. We use the standard cross-entropy loss instead of BCE, since the unnormalized language head weights could otherwise grow arbitrarily large to minimize the loss. To convert GPT-2 and Llama-3.2 to B-cos LMs, we apply B-cos adaptations and further train them on 500,000 and 4,000,000 sentences from OpenWebText [^4] https://huggingface.co/datasets/Skylion007/openwebtext, respectively.
Datasets
For explanation evaluation, we use the BLiMP dataset (Warstadt et al., 2020) to assess explanations for linguistic phenomena, and the Indirect Object Identification (IOI) dataset (Brian Muhia, 2022) to test models’ reasoning about object identification. Following Ferrando et al. (2023), we use nine subsets of BLiMP. Each example in both datasets consists of a sentence prefix followed by a target and a foil next-word prediction, differing in whether they align with the phenomenon or ability of interest. Ground truth evidence is provided to support either grammatical correctness or correct object identification. Examples of these datasets can be found in Table 8 in Appendix C.
Metrics and Baselines
We evaluate explanation quality using Mean Reciprocal Rank (MRR), where higher scores indicate stronger alignment with the ground truth evidence. To assess language modeling abilities of models, we report two metrics: (1) the probability gap between target and foil predictions, and (2) perplexity (PPL) on a held-out corpus. Following Yin & Neubig (2022), we generate contrastive explanations that explain why the model predicts target tokens instead of foil tokens, and compare B-cos explanations against several baseline methods: L1 gradient norm (Grad Norm), IxG, Occlusion, and two propagation-based methods Logit and ALTI Logit from Ferrando et al. (2023).
Method GPT-2 MRR ($\uparrow$) Llama-3.2 MRR ($\uparrow$)
BLiMP IOI BLiMP IOI
Random 0.5130 0.2360 0.5132 0.2328
Grad Norm 0.5465 0.8599 0.5504 0.3637
IxG 0.4750 0.1112 0.5303 0.1034
Occlusion 0.6365 0.8517 0.6201 0.4767
Logit 0.7307 1.0 - -
ALTI Logit 0.7391 1.0 - -
B-cos 0.7561 1.0 0.6969 0.9913
Table 7: Alignment results (MRR) on BLiMP and IOI. BLiMP MRR scores are averaged over nine subsets, and Logit and ALTI Logit results are replicated from the original paper (Ferrando et al., 2023). Best scores are marked in bold. B-cos explanations achieve the best alignment with ground truth evidence. Results for each subset can be found in Table 14 and Table 15 in Appendix N.
Results
Table 6 shows that B-cos GPT-2 and B-cos Llama-3.2 models achieve strong language modeling performance comparable to their vanilla counterparts. Besides, Table 7 demonstrates that B-cos explanations exhibit better alignment with ground truth across tasks and models, indicating improved explainability of B-cos decoder-only LMs. Although the current B-cosification pipeline requires additional training, future work could explore more efficient approaches that reduce training overhead or integrate B-cosification into the pre-training phase. Overall, we believe B-cos decoder-only models are well-suited for tasks where explainability is critical and represent a promising direction for building more transparent and reliable LLMs.
@@ 8 Conclusion
In this work, we introduce B-cos LM, a bias-free dynamic linear model that learns task-relevant patterns through increased input-weight alignment pressure. B-cos LMs generate more faithful and human interpretable explanations while maintaining strong task performance and fast convergence. Based on our in-depth analysis of B-cosification, we provide three recommendations for effectively transforming PLMs into B-cos LMs: (1) combine B-cos conversion and task fine-tuning for efficient B-cosification. If resources allow, additional B-cos pre-training can further improve task performance and explanation faithfulness; (2) carefully select the parameter B, as excessively large values can reduce model capacity and lead to overly sparse explanations; and (3) be mindful of biases in training data, especially at high B values, as B-cosification may amplify existing biases.
We also explore adapting decoder-only models into B-cos LMs for generation tasks and show that, with additional training, they match the language modeling performance of conventional models while providing better explanations. We hope these findings support future efforts in building explainable LLMs.
@@ 9 Limitations
Second, we find that B-cos explanations do not consistently capture token interactions within multi-token phrases. For example, a negation phrase like not good tends to receive an overall attribution score that aligns with its meaning (e.g., a negative score for positive sentiment), but the individual token scores within the phrase vary across contexts. In some cases, the word good may receive either positive or negative scores across different examples, even when the overall sentiment remains the same. Similar issues arise in other methods, suggesting a broader limitation of token-level rationales in capturing compositional semantics.
10 Broader Impact Statement
@@ Appendix A Terminology
  • Faithfulness The extent to which an explanation accurately reflects the model’s actual reasoning process (Jacovi & Goldberg, 2020). A faithful explanation should directly correspond to the internal mechanisms that led to the model’s prediction.
  • Human Interpretability The ease with which a person can understand the model’s reasoning from the explanation (Lage et al., 2019). A highly interpretable explanation should be clear, concise, and focused on relevant information while avoiding unnecessary or distracting information. However, an explanation that is easy for humans to interpret may not necessarily reflect the model’s actual reasoning process or align with human reasoning patterns.
  • Human Agreement The degree to which a model’s explanation aligns with the reasoning a human would use for the same prediction. A high-agreement explanation should follow intuitive, logical reasoning patterns similar to human decision-making.
  • Explainability The extent to which a model’s computations can be faithfully explained and its learned patterns are understandable to humans. A highly explainable model should yield explanations that are both faithful to its actual reasoning process and interpretable to humans.
@@ Appendix B Notation
In this paper, we use lowercase letters for scalars (e.g., b), bold lowercase letters for vectors (e.g., $\mathbf{w}$, $\mathbf{x}$), and bold uppercase letters ($\mathbf{W}$) for matrices.A special case is the alignment pressure parameter, denoted by the non-bold uppercase letter B, to distinguish it from the bias term b in linear layers. We use bold uppercase letters $\mathbf{X}$ and $\mathbf{A}$ to denote a sequence of model inputs or hidden state activations. In § 3, we use $\mathbf{x}$ to denote the input when a function is applied to each element of the input sequence separately. In contrast, we use $\mathbf{X}$ or $\mathbf{A}$ when the function involves interactions between elements, such as in the attention mechanism.
Appendix C Implementation Details
@@
For Saloss models, we use the official codebase [^9] https://github.com/GChrysostomou/saloss with default hyperparameters to train BERT and RoBERTa on AG News, IMDB, and HateXplain. DistilBERT is not included, as it is not supported by the codebase.
In Section 7, we follow Ferrando et al. (2023) to generate contrastive explanations that highlight why the models predicts the target token instead of the foil token. For Occlusion explanations, we use the [PAD] token to perform occlusion, instead of a zero vector as done in Yin & Neubig (2022) and Ferrando et al. (2023). Using zero vectors distorts the input distribution and, in generative settings, can influence predictions differently depending on position. To avoid such positional effects, we instead occlude using the in-distribution embedding of the [PAD] token.
@@
When constructing examples for SeqPG, we set the sequence length to 50 for AG News, 256 for IMDB, and 25 for HateXplain, aligning with their median lengths. Only examples longer than these thresholds are selected, and they are truncated to construct synthetic examples. Additionally, we only use examples that are correctly predicted with a minimum confidence of 75% after truncation. For a fair comparison, we evaluate Saloss models and B-cos LMs on the same sets of examples constructed based on the predictions of the corresponding conventional models.
Automatic Evaluation Setups
For task performance evaluation, we use the complete test set for each task. For faithfulness evaluation, we conduct perturbation-based evaluations on 2000 test examples and SeqPG on 500 test examples for AG News and IMDB. For HateXplain, we use the full test set for perturbation-based evaluation (1,924 examples) and construct 269, 310, and 308 SeqPG examples from it using BERT, DistilBERT, and RoBERTa, respectively. In the perturbation-based evaluation, the [CLS] token is never perturbed because it is used directly to make predictions.
@@
For B-cos pre-training in § 5, we set B=1.25 and further pre-train the model on 25,000 sentences from the Wikipedia dataset [^10] https://huggingface.co/datasets/wikimedia/wikipedia using masked language modeling loss with a learning rate of 1e-4 and a 15% masking ratio. We do not B-cosify the language head, as its parameters are tied with the embedding layer. Pre-training uses the standard cross-entropy loss rather than binary cross-entropy loss, since the unnormalized language head weights could otherwise grow arbitrarily large to minimize the loss.
Decoder-only Models B-cosification
We use the GPT-2 small and Llama-3.2-1B models from huggingface. As with the encoder-based models, we do not B-cosify the language head and use cross-entropy loss for GPT-2 and Llama-3.2 training. B-cos adaptations are first applied and both models are then trained on 500,000 and 4,000,000 sentences, respectively, from the OpenWebText dataset for one epoch, using a learning rate of 5e-4. For GPT-2, we use a batch size of 16 and a maximum sequence length of 512; for Llama-3.2, we use a batch size of 128 and a sequence length of 1024. Perplexity is evaluated on a held-out OpenWebText subset of 10,000 sentences using a maximum sequence length of 512.
BLiMP Subsets
We follow Ferrando et al. (2023) to use the following nine BLiMP subsets with corresponding IDs. aga: anaphor_gender_agreement; ana: anaphor_number_agreement; asp: animate_subject_passive; dna: determiner_noun_agreement_1; dnai: determiner_noun_agreement_irregular_1; dnaa: determiner_noun_agreement_with_adj_1; dnaai: determiner_noun_agreement_with_adj_irregular_1; npi: npi_present_1; darn: distractor_agreement_relational_noun. Examples of these datasets and the IOI dataset can be found in Table 8.
Dataset ID Example
Anaphor gender agreement aga Katherine can’t help herself / himself.
Anaphor number agreement ana Susan revealed herself / themselves.
Animate subject passive asp Amanda was respected by some waitresses / pictures.
Determiner noun agreement 1 dna Raymond is selling this sketch / sketches.
Determiner noun agreement irregular 1 dnai Adam hadn’t discussed these analyses / analysis.
Determiner noun agreement with adjectives 1 dnaa Rebecca was criticizing those good documentaries / documentary.
Determiner noun agreement with adjectives irregular 1 dnaai Some waiters broke this lost foot / feet.
NPI present 1 npi Even Suzanne has really / ever joked around.
Distractor agreement relational noun darn A niece of most senators hasn’t / haven’t descended most slopes.
Indirect Object Identification IOI Friends Juana and Kristi found a mango at the bar.
Kristi gave it to Juana / Kristi.
Table 8: Examples from the BLiMP and IOI datasets. Green (red) indicates target (foil) predictions. Ground truth evidence for the correct continuations is underlined.
@@
Appendix D SeqPG Example